# Axioms for AI Alignment from Human Feedback

### v1

@@ 1 Introduction
The alignment of AI models with human values is widely recognized as a crucial task. A prominent method for this task, reinforcement learning with human feedback (RLHF), has been used in different applications, such as robotics Bıyık et al. (2020); Kupcsik et al. (2018) and recommendations Viappiani and Boutilier (2010); Ailon and Mohri (2010). Recently, RLHF has attracted significant attention as a tool for fine-tuning large language models (LLMs) Ouyang et al. (2022); Ziegler et al. (2019); Stiennon et al. (2020). A typical implementation of RLHF involves learning a reward model using a pre-trained LLM, which is then utilized to fine-tune an existing LLM. During the learning step, human feedback is provided in the form of ordinal comparisons, and a reward function is learned from these. The most common learning method assumes an underlying random utility model such as the Bradley-Terry-Luce (BTL) model Ouyang et al. (2022); Christiano et al. (2017) and computes a reward function that corresponds to a maximum likelihood estimator for the observed comparisons.
Is this the “right” way of aggregating individual preferences towards a socially desirable reward function? To answer this question, we draw on social choice theory, a field that studies collective decison making through a mathematical lens Brandt et al. (2016). The maximum likelihood estimation approach is in line with a storied body of work that assumes that different human participants have preferences stemming from noisy estimation of a common ground truth, and the goal is to learn this ground truth as accurately as possible Young (1988). But this is not the case when it comes to questions of AI alignment, where individuals can have legitimate differences of opinion rooted in different values or priorities.
We argue that when preferences are truly heterogeneous, the axiomatic approach — which rose to prominence in social choice with the work of Arrow (1951) — may be more suitable. This approach analyzes the desirability of aggregation methods through their satisfaction of certain axioms that capture notions of consensus, fairness, and economic efficiency. Specifically, we are interested in the axiomatic properties of aggregation methods that take ordinal preferences as input and output a reward function. We address the following two research questions: What axioms are satisfied by aggregation methods used by existing RLHF algorithms? And are there alternative aggregation methods that offer stronger axiomatic guarantees?
@@ 1.1 Our Approach
In social choice theory, axioms are typically defined for rules that map rankings over candidates to a single winner (social choice functions) or a ranking of the candidates (social welfare functions). By contrast, we are interested in rules that assign a reward to each candidate. This gap is easy to bridge, though: we simply consider a ranking of the candidates by decreasing reward.
@@
Specifically, we assume that a linear reward function defined by a parameter vector determines the reward of each candidate by computing the inner product of the parameter vector and the feature vector of the candidate; these modeling choices are consistent with prior and concurrent work Zhu et al. (2023); Zhong et al. (2024); Ge et al. (2024) and aim to capture the practice of RLHF. [^1] We can represent the reward model as an embedding layer $\phi(x)$ which is then linearly aggregated to compute the final reward. If we fix the embedding function $\phi$ and treat its output as the feature representation of the outcomes, such as prompt-response pairs, the resulting reward model is linear in $\phi(x)$; see, e.g. Zhu et al. (2023). Each human participant (henceforth referred to as a voter) is associated with a parameter vector, which is unknown to us and is used to specify ordinal preferences over the candidates. Our task is to design linear rank aggregation rules, which aggregate rankings induced by these individual linear functions [^2] In practice, typical RLHF datasets consist of pairwise comparisons, not complete rankings. Assuming rankings as input makes our exposition cleaner, and it is not a fundamental limitation, as we discuss in Section 5. into a collective ranking that is also induced by a linear function; this is a new paradigm in social choice, for which we coin the term linear social choice.
To evaluate linear rank aggregation rules, we adapt fundamental axioms from social choice theory. The first is Pareto optimality (PO), which requires that if a candidate $a$ is ranked above candidate $b$ in every input ranking, then the resulting ranking should rank $a$ above $b$. This is seen as a basic requirement and is satisfied by every standard voting method in the classical setting. The second axiom is pairwise majority consistency (PMC): If there exists a reward function that generates a ranking where, for each pair of candidates, a majority of voters agree with the ranking, then the resulting ranking should match that ranking. This axiom is an extension of Condorcet consistency to rankings, and is satisfied by some, but not all, standard voting methods in the classical setting.
@@ 1.2 Our Results
In Section 3.2, we take a first step towards addressing this shortcoming. We modify the loss-based formulation to focus on majority preferences instead of those of individuals. This modification defines a family of rules that are PMC, but we show that all of them fail PO by establishing an even stronger impossibility result: In stark contrast to the classical setting, any linear rank aggregation rule that depends only on majority preferences must fail PO.
@@ 2 The Linear Social Choice Model
Note that a PMC ranking for each $\pi$ need not exist, but when one does, it is unique. The words “$\sigma$ is feasible” allude to the possibility that no non-degenerate parameter vector $\theta$ induces the unique PMC ranking. Indeed, we have the such an example; see Appendix A for details.
@@ 3.1 Standard Loss Formulation
With this definition in hand, we proceed to our first main result, which spells rather bad news for this class of rules: Any loss-based aggregation rule using a nondecreasing and convex loss function (of which BTL is one, and hinge loss is another) will fail our two core axioms, PMC and PO. This paints a negative picture for current RLHF methods with respect to their social choice guarantees. Note that we will exclude the discussion of loss functions that have its global minimum at zero like Relu because the loss minimizer will be zero, making all rankings vacuous consequently. And we have focused on convex loss functions due to its practical ease in optimization.
@@
Let $OPT^{core}=\{\theta\mid\mathcal{L}^{core}(\theta)=\inf_{\theta^{\prime}}\mathcal{L}^{core}(\theta^{\prime})\}$ and $OPT^{unconstr}\{(r_{a},r_{b})\mid\mathcal{L}^{unconstr}(r_{a},r_{b})=\inf_{r^{\prime}_{a},r^{\prime}_{b}}\mathcal{L}^{unconstr}(r^{\prime}_{a},r^{\prime}_{b})\}$ be the set of minimizers for these two loss functions. We now establish the following results about these optimal sets.
@@
Proof of Lemma 3.2.
@@
We begin with some observations about $g$. First, we have that since $\ell$ is nonnegative, $g$ is nonnegative. This along with the fact that $\lim_{x\to\infty}\ell(x)=\infty$, we have that both $\lim_{x\to\infty}g(x)=\infty$ and $\lim_{x\to-\infty}g(x)=\infty$. Together, these imply that $\mathcal{L}^{unconstr}$ attains a minimum. Indeed, $\mathcal{L}^{unconstr}(0,0)=3g(0)$, and there is some bound $B$ such that for all $x>B$ and $x<-B$, $g(x)>3g(0)$. We can therefore restrict the optimization problem to $r_{a},r_{b}\in[-B,B]$ without changing the solutions. Since $\mathcal{L}^{unconstr}$ is continuous and $[-B,B]^{2}$ is compact, a minimum is attained.
@@
Next, note that $g$ is convex because compositions of convex functions with monotonic functions and convex combinations of convex functions are convex Rockafellar (1970). From this, we claim that if there is an optimal solution $(r_{a},r_{b})$, then $(r_{a},r_{a}/2)$ is also an optimal solution. Indeed, fix such an $(r_{a},r_{b})$,
@@
Further, if $k$ is convex, we have that (i) $k^{\prime}_{+}$ and $k^{\prime}_{-}$ are nondecreasing, (ii) for every $x$, $k^{\prime}_{-}(x)\leq k^{\prime}_{+}(x)$, (iii) $x$ minimizes $k$ if and only if $k^{\prime}_{-}(x)\leq 0\leq k^{\prime}_{+}(x)$, and (iv) $k^{\prime}_{+}$ and $k^{\prime}_{-}$ are right and left continuous, respectively Rockafellar (1970). They also follow standard linearity and chain rule properties, which allow for simpler computation, e.g., if $k(x)=a\alpha(x)+b\beta(x)$, then $k^{\prime}_{+}(x)=a\alpha^{\prime}_{+}(x)+b\beta^{\prime}_{+}(x)$ and if $k(x)=\alpha(\gamma x)$, then $k_{+}^{\prime}(x)=\gamma\alpha^{\prime}_{+}(\gamma x)$ if $\gamma\geq 0$, and $k_{+}^{\prime}(x)=\gamma\alpha^{\prime}_{0}(\gamma x)$ (all of these for left-hand derivatives swapping the places of $+$ and --Clarke (1983).
@@
First, suppose $\ell$ is non-decreasing. This implies that $\ell^{\prime}_{-}(x)\geq 0$ for all $x$. We know that there is a point $x<0$ such that $\ell^{\prime}_{-}(x)<\ell^{\prime}_{-}(0)$ (as otherwise $\ell$ would eventually become negative). Let $d=\ell^{\prime}_{-}(x)$. Take $w$ such that $-w=\sup\{x\mid\ell^{\prime}_{-}(x)=d\}$. Note that $\ell^{\prime}_{-}(-w)=d$ because $\ell^{\prime}_{-}$ is left continuous (from (iv) above). Since $-w<0$, so $-w/2>-w$. This implies that $\ell^{\prime}_{-}(-w/2)>d$, as otherwise $-w$ would not be the supremum of such points. In addition, we have that $d\geq 0$ because $\ell$ is nondecreasing.
@@
For this choice of $w$, we first claim that the preconditions of denominators being positive hold for (2) and (3). Indeed, let us write $z_{1},z_{2},z_{3},z_{4}$ for $\ell^{\prime}_{-}(-2w),\ell^{\prime}_{-}(-w),\ell^{\prime}_{+}(w),\ell^{\prime}_{+}(2w)$. We know that $z_{1}\leq z_{2}\leq z_{3}\leq z_{4}$ by properties of convexity, and since $\ell^{\prime}_{-}(-w/2)>\ell^{\prime}_{-}(-w)\geq 0$, we have that $0\leq z_{1}<z_{2}$. The denominators are of the form $z_{1}+z_{4}$ and $z_{1}+z_{2}+z_{3}+z_{4}$, which are now both necessarily positive. In addition, we also claim that a rational $p$ with $11/2<p<$ 1 satisfying both inequalities (2) and (3) will exist. Note that inequality (2) can now be represented as $\frac{z_{4}}{z_{1}+z_{4}}$, and we have that $\frac{z_{4}}{z_{1}+z_{4}}\leq 1$ because $0\leq z_{1}<z_{4}$ Additionally, inequality (3) can be represented as $\frac{z_{3}+z_{4}}{z_{1}+z_{2}+z_{3}+z_{4}}$ which is at least $1/2$ because $z_{3}+z_{4}>z_{1}+z_{2}$. Finally, we have
@@
Since $M=\begin{pmatrix}2&1\\ 1&1\end{pmatrix}$ is invertible with inverse $M^{-1}=\begin{pmatrix}1&-1\\ -1&2\end{pmatrix}$, for any $r_{a},r_{b}$, we can simply set $\theta=M^{-1}\begin{pmatrix}r_{a}\\ r_{b}\end{pmatrix}$. Thus, $OPT^{com}$ is nonempty, and is simply the image of $OPT^{unconstr}$ under $M^{-1}$. Now, fix $(r_{a},r_{b})\in OPT^{unconstr}$, and let $\theta=M^{-1}\begin{pmatrix}r_{a}\\ r_{b}\end{pmatrix}$. By assumption, we have $r_{a}>A_{2}$ and $r_{b}\leq A_{1}$. This implies that $\theta_{1}=r_{a}-r_{b}>A_{2}-A_{1}$, while $\theta_{2}=2r_{b}-r_{a}<2A_{1}-A_{2}$. Thus, setting $A_{3}=A_{2}-A_{1}$ and $A_{4}=2A_{1}-A_{2}$ satisfy the desired properties. ∎
For each $\varepsilon\in\mathbb{R}$, let
Let $OPT(\varepsilon)$ be the set of vectors optimizing $\mathcal{L}^{\varepsilon}$. The rest of the proof will follow from the following two lemmas.
Lemma 3.4.
$OPT(0)\subseteq\overline{R^{c^{\prime}\succ c}}$.
Lemma 3.5.
Proof of Lemma 3.4.
@@
This implies that we may restrict our attention to $\theta$ in the region $R^{bounded}=\{\theta\mid|\theta_{1}|\leq B,|\theta_{2}|\leq 2B\}$. Indeed, for $\theta\notin R^{bounded}$, either $|\theta_{1}|>B$ or $|\theta_{1}+\theta_{2}|>B$. In either case, we have $\mathcal{L}^{\varepsilon}(\theta)\geq(1-p)\ell(B)>\mathcal{L}^{\varepsilon}(\mathbf{0})$.
Note that $\mathcal{L}^{\varepsilon}(\theta)$ is continuous not only in $\theta$, but also in $\varepsilon$. Additionally, $R^{bounded}$ is closed and bounded, and hence, compact. Therefore, by Berge’s Maximum Theorem, $OPT(\varepsilon)$ is nonempty and upper semi-continuous in $\varepsilon$ Berge (1963). As per the definition of upper semi-continuouus, since $OPT(0)\subseteq\overline{R^{c^{\prime}\succ c}}$, an open set, for sufficiently small $\varepsilon>0$, $OPT(\varepsilon)\subseteq\overline{R^{c^{\prime}\succ c}}$. Finally, note that $R^{c^{\prime}\succ c}\cap R^{bounded}$ is compact, so a minimum is attained, and this minimum must therefore be strictly larger than the values attained by members of $OPT(\varepsilon)$. ∎
3.2 Majority-Based Loss Formulation
We first show that this does indeed help achieve PMC with essentially all loss functions.
Theorem 3.6.
@@ 3.2 Majority-Based Loss Formulation
However, despite this good news for PMC, we show that this does not help in achieving PO. In fact, our negative result extends to every $C1$ linear rank aggregation rule. Note that if $f$ minimizing $\ell$ in the majority formulation breaks ties consistently (i.e., if multiple feasible rankings are optimal, then it consistently chooses the same one), then it is C1. We then have the following result.
Theorem 3.7.
All $C1$ linear rank aggregation rules fail PO.
@@
4 Social Choice-Based Rule
Copeland’s rule is a $C1$ rule because it only requires the majority relationships between the candidates. Analogously, leximax Copeland is also a $C1$ linear rank aggregation rule. Therefore, by Theorem 3.7, it does not satisfy the PO criterion. To address this issue, we define a variant called leximax Copeland subject to PO (LCPO), which incorporates the PO criterion. Under LCPO, for every pair of alternatives where one dominates the other, the rule restricts rankings to place the dominating alternative above the dominated one. The rule remains well-defined since the set of feasible rankings when enforcing the PO criterion is non-empty, as whenever $a$ dominates $b$, all the rankings in the input profile rank $a$ above $b$. Note that if the Copeland ranking is feasible, then this rule outputs that ranking, since unrestricted Copeland satisfies PO.
Definition 4.1 (majority consistency).
Definition 4.2 (winner monotonicity).
We now state and prove the main result of this section.
Theorem 4.3.
Proof.
5 Discussion
Second, as written, our model has voters give their complete rankings, while in practice, this would be infeasible. In the real world, we are likely to elicit only relatively few pairwise comparisons per person. For our negative results, this assumption only makes them stronger: the BTL model fails PO and PMC even when it has access to complete voter rankings. By contrast, for the positive results, specifically implementing leximax Copeland subject to PO, this ostensibly seems like a serious limitation. However, the complete rankings are not necessary for computing this rule, rather, all we need to know are PO dominance relationships and majority directions. We can therefore apply the rule whenever we can determine this information at least approximately through, e.g., sampling. An alternative approach is to infer a complete ranking of each voter by fitting a parameter vector based on their pairwise responses; this process of learning a complete ranking and then running voting rules has been used before in a variety of settings Noothigattu et al. (2018); Lee et al. (2019)
Some rules which optimize over rankings can be naturally transformed. For example, consider the Kemeny rule, which returns the ranking with the smallest pairwise disagreement over all votes. This can easily be transformed to the linear setting by simply outputting the optimal feasible ranking. In fact, in Section B.2, we show that this rule carries over the property of separability, [^9] Formally, ranking separability to distinguish it from the single-winner version. a social choice axiom that is violated by Copeland (in the classical setting) and leximax Copeland subject to PO (in our setting). We show this in Section B.1. However, quite strikingly, although separability remains, this transformation makes Kemeny no longer PO (Section B.2).
Finally, note that the “leximax” portion of leximax Copeland can be be seen as a general purpose tool for mapping traditional rules to a linear aggregation ones. In Section B.3, we explore leximax plurality (run leximax on the ranking of candidates by plurality scores), and show that it satisfies majority consistency, winner monotonicity, and separability. Additionally, the “subject to PO” can be seen as another “tool” for enforcing the Pareto optimality criterion when a rule does not independently satisfy it. However, enforcing PO can again cause somewhat surprising results. For example, in Section B.2, we show that linear Kemeny subject to PO, while now trivially satisfying PO, again violates separability. These observations indicate the challenges inherent in linear social choice, and we hope the many open questions will lead to fruitful followup research.
Appendix A PMC Infeasibility Example

### v2

@@ 1 Introduction
The alignment of AI models with human values is widely recognized as a crucial task. A prominent method for this task, reinforcement learning with human feedback (RLHF), has been used in different applications, such as robotics Bıyık et al. (2020); Kupcsik et al. (2018) and recommendations Viappiani and Boutilier (2010); Ailon and Mohri (2010). Recently, RLHF has attracted significant attention as a tool for fine-tuning large language models (LLMs) Ouyang et al. (2022); Ziegler et al. (2019); Stiennon et al. (2020). A typical implementation of RLHF involves learning a reward model using a pre-trained LLM, which is then utilized to fine-tune an existing LLM. During the learning step, human feedback is provided in the form of ordinal comparisons, and a reward function is learned from these. The most common learning method assumes an underlying random utility model such as the Bradley-Terry-Luce (BTL) model Bradley and Terry (1952); Ouyang et al. (2022); Christiano et al. (2017) and computes a reward function that corresponds to a maximum likelihood estimator for the observed comparisons.
Is this the “right” way of aggregating individual preferences towards a socially desirable reward function? To answer this question, we draw on social choice theory, a field that studies collective decision making through a mathematical lens Brandt et al. (2016). The maximum likelihood estimation approach is in line with a well-established body of work that assumes that different human participants have preferences stemming from noisy estimation of a common ground truth, and the goal is to learn this ground truth as accurately as possible Young (1988). But this is not the case when it comes to questions of AI alignment, where individuals can have legitimate differences of opinion rooted in different values or priorities.
We argue that when preferences are truly heterogeneous, the axiomatic approach — which rose to prominence in social choice with the work of Arrow (1951) — may be more suitable. This approach analyzes the desirability of aggregation methods by their satisfaction of certain axioms that capture notions of consensus, fairness, and economic efficiency. Specifically, we are interested in the axiomatic properties of aggregation methods that take ordinal preferences as input and output a reward function. We address the following two research questions: What axioms are satisfied by aggregation methods used by existing RLHF algorithms? And are there alternative aggregation methods that offer stronger axiomatic guarantees?
@@ 1.1 Our Approach
In social choice theory, axioms are typically defined for rules that map rankings over candidates to a single winner (social choice functions) or a ranking of the candidates (social welfare functions). By contrast, we are interested in rules that assign a reward to each candidate. This gap is easy to bridge, though: we simply consider a ranking of the candidates by in descending reward order.
@@
Specifically, we assume that a linear reward function defined by a parameter vector determines the reward of each candidate by computing the inner product of the parameter vector and the feature vector of the candidate; these modeling choices are consistent with prior and concurrent work Zhu et al. (2023); Zhong et al. (2024); Ge et al. (2024) and aim to capture the practice of RLHF. [^1] We can represent the reward model as an embedding layer $\phi(x)$ which is then aggregated linearly to compute the final reward. If we fix the embedding function $\phi$ and treat its output as the feature representation of the outcomes, such as prompt-response pairs, the resulting reward model is linear in $\phi(x)$; see, e.g. Zhu et al. (2023). Each human participant (henceforth referred to as a voter) is associated with a parameter vector, which is unknown to us and is used to specify ordinal preferences over the candidates. Our task is to design linear rank aggregation rules, which aggregate rankings induced by these individual linear functions [^2] In practice, typical RLHF datasets consist of pairwise comparisons, not complete rankings. Assuming rankings as input makes our exposition cleaner, and it is not a fundamental limitation, as we discuss in Section 5. into a collective ranking that is also induced by a linear function; this is a new paradigm in social choice, for which we coin the term linear social choice.
To evaluate linear rank aggregation rules, we adapt fundamental axioms from social choice theory Brandt et al. (2016). The first is Pareto optimality (PO), which requires that if a candidate $a$ is ranked above candidate $b$ in every input ranking, then the resulting ranking should rank $a$ above $b$. This is seen as a basic requirement and is satisfied by every standard voting method in the classical setting.
The second axiom is pairwise majority consistency (PMC): If there exists a reward function that generates a ranking where, for each pair of candidates, a majority of voters agree with the ranking, then the resulting ranking should match that ranking. This axiom is an extension of Condorcet consistency to rankings, and is satisfied by some, but not all, standard voting methods in the classical setting.
@@ 1.2 Our Results
In Section 3.2, we take a first step towards addressing this shortcoming. We modify the loss-based formulation to focus on majority preferences rather than individual preferences. This modification defines a family of rules that are PMC, but we show that all of them fail PO by establishing an even stronger impossibility result: In stark contrast to the classical setting, any linear rank aggregation rule that depends only on majority preferences must fail PO.
@@ 2 The Linear Social Choice Model
Note that a PMC ranking for each $\pi$ need not exist, but when one does, it is unique. The words “$\sigma$ is feasible” allude to the possibility that no non-degenerate parameter vector $\theta$ induces the unique PMC ranking. Indeed, we have such an example; see Appendix B for details.
@@ 3.1 Standard Loss Formulation
With this definition in hand, we proceed to our first main result, which spells rather bad news for this class of rules: Any loss-based aggregation rule using a nondecreasing and convex loss function (of which BTL is one, and hinge loss is another) will fail our two core axioms, PMC and PO. This paints a negative picture for current RLHF methods with respect to their social choice guarantees. Note that we will exclude the discussion of loss functions with a global minimum at zero, like ReLU, because the loss minimizer will be zero, making all rankings vacuous consequently. And we have focused on convex loss functions due to their practical optimization ease.
@@
Let $OPT^{core}=\{\theta\mid\mathcal{L}^{core}(\theta)=\inf_{\theta^{\prime}}\mathcal{L}^{core}(\theta^{\prime})\}$ and $OPT^{unconstr}\{(r_{a},r_{b})\mid\mathcal{L}^{unconstr}(r_{a},r_{b})=\inf_{r^{\prime}_{a},r^{\prime}_{b}}\mathcal{L}^{unconstr}(r^{\prime}_{a},r^{\prime}_{b})\}$ be the set of minimizers for these two loss functions. In Appendix A, we establish the following results about these optimal sets.
@@
For each $\varepsilon\in\mathbb{R}$, let
Let $OPT(\varepsilon)$ be the set of vectors optimizing $\mathcal{L}^{\varepsilon}$. The rest of the proof will follow from the following two lemmas, whose proofs are in Appendix A.
Lemma 3.4.
$OPT(0)\subseteq\overline{R^{c^{\prime}\succ c}}$.
Lemma 3.5.
3.2 Majority-Based Loss Formulation
We first show (in Section A.5) that this does indeed help achieve PMC with essentially all loss functions.
Theorem 3.6.
However, despite this good news for PMC, we show that this does not help in achieving PO. In fact, our negative result extends to every $C1$ linear rank aggregation rule. Note that if $f$ minimizing $\ell$ in the majority formulation breaks ties consistently (i.e., if multiple feasible rankings are optimal, then it consistently chooses the same one), then it is C1. We then have the following result, whose proof is relegated to Section A.6.
Theorem 3.7.
All $C1$ linear rank aggregation rules fail PO.
4 Social Choice-Based Rule
Copeland’s rule is a $C1$ rule because it only requires the majority relationships between the candidates. Analogously, leximax Copeland is also a $C1$ linear rank aggregation rule. Therefore, by Theorem 3.7, it does not satisfy the PO criterion. To address this issue, we define a variant called leximax Copeland subject to PO (LCPO), which incorporates the PO criterion. Under LCPO, for every pair of alternatives where one dominates the other, the rule restricts rankings to place the dominating alternative above the dominated one.
The rule remains well-defined since the set of feasible rankings when enforcing the PO criterion is non-empty, as whenever $a$ dominates $b$, all the rankings in the input profile rank $a$ above $b$. Note that if the Copeland ranking is feasible, then this rule outputs that ranking, since unrestricted Copeland satisfies PO.
Definition 4.1 (majority consistency).
Definition 4.2 (winner monotonicity).
We now state and prove the main result of this section.
Theorem 4.3.
Proof.
5 Discussion
Second, as written, our model has voters give their complete rankings, while in practice, this would be infeasible. In the real world, we are likely to elicit only relatively few pairwise comparisons per person. For our negative results, this assumption only makes them stronger: the BTL model fails both PO and PMC even with access to complete voter rankings. By contrast, for the positive results, specifically implementing leximax Copeland subject to PO, this ostensibly seems like a serious limitation. However, the complete rankings are not necessary for computing this rule, rather, all we need to know are PO dominance relationships and majority directions. We can therefore apply the rule whenever we can approximate this information, for example, through sampling. An alternative approach is to infer a complete ranking of each voter by fitting a parameter vector based on their pairwise responses; this process of learning a complete ranking and then running voting rules has been used before in a variety of settings Noothigattu et al. (2018); Lee et al. (2019).
Some rules which optimize over rankings can be naturally transformed. For example, consider the Kemeny rule, which returns the ranking with the smallest pairwise disagreement over all votes. This can easily be transformed to the linear setting by simply outputting the optimal feasible ranking. In fact, in Section C.2, we show that this rule carries over the property of separability, [^9] Formally, ranking separability to distinguish it from the single-winner version. a social choice axiom that is violated by Copeland (in the classical setting) and leximax Copeland subject to PO (in our setting). We show this in Section C.1. However, quite strikingly, although separability remains, this transformation makes Kemeny no longer PO (Section C.2).
Finally, note that the “leximax” portion of leximax Copeland can be seen as a general purpose tool for mapping traditional rules to linear aggregation rules. In Section C.3, we explore leximax plurality (run leximax on the ranking of candidates by plurality scores), and show that it satisfies majority consistency, winner monotonicity, and separability. Additionally, the “subject to PO” can be seen as another “tool” for enforcing the Pareto optimality criterion when a rule does not independently satisfy it. However, enforcing PO can again cause somewhat surprising results. For example, in Section C.2, we show that linear Kemeny subject to PO, while now trivially satisfying PO, again violates separability. These observations indicate the challenges inherent in linear social choice, and we hope these open questions inspire fruitful follow-up research.
Acknowledgments
This research was partially supported by the National Science Foundation under grants IIS-2147187, IIS-2229881, CCF-2007080, IIS-1905558, and IIS-2214141; and by the Office of Naval Research under grants N00014-20-1-2488 and N00014-24-1-2704.
Appendix A Deferred Proofs
A.1 Proof of Lemma 3.2
Proof.
@@ Appendix A Deferred Proofs
We begin with some observations on $g$. First, we have that since $\ell$ is nonnegative, $g$ must also be nonnegative. This along with the fact that $\lim_{x\to\infty}\ell(x)=\infty$, we have that both $\lim_{x\to\infty}g(x)=\infty$ and $\lim_{x\to-\infty}g(x)=\infty$. Together, these imply that $\mathcal{L}^{unconstr}$ attains a minimum. Indeed, $\mathcal{L}^{unconstr}(0,0)=3g(0)$, and there is some bound $B$ such that for all $x>B$ and $x<-B$, $g(x)>3g(0)$. We can therefore restrict the optimization problem to $r_{a},r_{b}\in[-B,B]$ without changing the solutions. Since $\mathcal{L}^{unconstr}$ is continuous and $[-B,B]^{2}$ is compact, a minimum is attained.
@@
Next, note that $g$ is convex because compositions of convex functions with monotonic functions and convex combinations of convex functions are convex Rockafellar [1970]. From this, we claim that if there is an optimal solution $(r_{a},r_{b})$, then $(r_{a},r_{a}/2)$ is also an optimal solution. Indeed, fix such an $(r_{a},r_{b})$,
@@
Further, if $k$ is convex, we have that Rockafellar [1970]:
[…]
They also follow standard linearity and chain rule properties, which allow for simpler computation. For example:
$\text{if }k(x)=a\alpha(x)+b\beta(x),\text{ then }k^{\prime}_{+}(x)=a\alpha^{\prime}_{+}(x)+b\beta^{\prime}_{+}(x),$
$\text{if }k(x)=\alpha(\gamma x),\text{ then }k^{\prime}_{+}(x)=\gamma\alpha^{\prime}_{+}(\gamma x)\text{ if }\gamma\geq 0,\text{ and }k^{\prime}_{+}(x)=\gamma\alpha^{\prime}_{-}(\gamma x)\text{ if }\gamma<0,$
(all of these hold for left-hand derivatives by swapping the positions of $+$ and --Clarke [1983].
@@
First, suppose $\ell$ is non-decreasing. This implies that $\ell^{\prime}_{-}(x)\geq 0$ for all $x$. We know that there is a point $x<0$ such that $\ell^{\prime}_{-}(x)<\ell^{\prime}_{-}(0)$ (as otherwise $\ell$ would eventually become negative). Let $d=\ell^{\prime}_{-}(x)$. Take $w$ such that
$-w=\sup\{x\mid\ell^{\prime}_{-}(x)=d\}.$
Note that $\ell^{\prime}_{-}(-w)=d$ because $\ell^{\prime}_{-}$ is left continuous (from (iv) above). Since $-w<0$, so $-\frac{w}{2}>-w$. This implies that $\ell^{\prime}_{-}(-\frac{w}{2})>d$, as otherwise $-w$ would not be the supremum of such points. In addition, we have that $d\geq 0$ because $\ell$ is nondecreasing.
@@
For this choice of $w$, we first claim that the preconditions of denominators being positive hold for (2) and (3). Indeed, let us write $z_{1},z_{2},z_{3},z_{4}$ for $\ell^{\prime}_{-}(-2w),\ell^{\prime}_{-}(-w),\ell^{\prime}_{+}(w),\ell^{\prime}_{+}(2w)$. We know that
$z_{1}\leq z_{2}\leq z_{3}\leq z_{4}$
by properties of convexity, and since $\ell^{\prime}_{-}(-w/2)>\ell^{\prime}_{-}(-w)\geq 0$, we have that $0\leq z_{1}<z_{2}$. The denominators are of the form $z_{1}+z_{4}$ and $z_{1}+z_{2}+z_{3}+z_{4}$, which are now both necessarily positive. In addition, we also claim that a rational $p$ with $11/2<p<$ 1 satisfying both inequalities (2) and (3) will exist. Note that inequality (2) can now be represented as $\frac{z_{4}}{z_{1}+z_{4}}$, and we have that
$\frac{z_{4}}{z_{1}+z_{4}}\leq 1$
because $0\leq z_{1}<z_{4}$. Additionally, inequality (3) can be represented as
$\frac{z_{3}+z_{4}}{z_{1}+z_{2}+z_{3}+z_{4}}$
which is at least $1/2$ because $z_{3}+z_{4}>z_{1}+z_{2}$. Finally, we have
@@
Since $M=\begin{pmatrix}2&1\\ 1&1\end{pmatrix}$ is invertible with inverse
$M^{-1}=\begin{pmatrix}1&-1\\ -1&2\end{pmatrix},$
for any $r_{a},r_{b}$, we can simply set
$\theta=M^{-1}\begin{pmatrix}r_{a}\\ r_{b}\end{pmatrix}.$
Thus, $OPT^{com}$ is nonempty, and is simply the image of $OPT^{unconstr}$ under $M^{-1}$. Now, fix $(r_{a},r_{b})\in OPT^{unconstr}$, and let $\theta=M^{-1}\begin{pmatrix}r_{a}\\ r_{b}\end{pmatrix}$. By assumption, we have $r_{a}>A_{2}$ and $r_{b}\leq A_{1}$. This implies that $\theta_{1}=r_{a}-r_{b}>A_{2}-A_{1}$, while $\theta_{2}=2r_{b}-r_{a}<2A_{1}-A_{2}$. Thus, setting $A_{3}=A_{2}-A_{1}$ and $A_{4}=2A_{1}-A_{2}$ satisfy the desired properties. ∎
A.3 Proof of Lemma 3.4
Proof.
@@
This implies that we may restrict our attention to $\theta$ in the region
$R^{bounded}=\{\theta\mid|\theta_{1}|\leq B,|\theta_{2}|\leq 2B\}.$
Indeed, for $\theta\notin R^{bounded}$, either $|\theta_{1}|>B$ or $|\theta_{1}+\theta_{2}|>B$. In either case, we have $\mathcal{L}^{\varepsilon}(\theta)\geq(1-p)\ell(B)>\mathcal{L}^{\varepsilon}(\mathbf{0})$.
Note that $\mathcal{L}^{\varepsilon}(\theta)$ is continuous not only in $\theta$, but also in $\varepsilon$. Additionally, $R^{bounded}$ is closed and bounded, and hence, compact. Therefore, by Berge’s Maximum Theorem, $OPT(\varepsilon)$ is nonempty and upper semi-continuous in $\varepsilon$ Berge [1963]. As per the definition of upper semi-continuous, since $OPT(0)\subseteq\overline{R^{c^{\prime}\succ c}}$, an open set, for sufficiently small $\varepsilon>0$, $OPT(\varepsilon)\subseteq\overline{R^{c^{\prime}\succ c}}$. Finally, note that $R^{c^{\prime}\succ c}\cap R^{bounded}$ is compact, so a minimum is attained, and this minimum must therefore be strictly larger than the values attained by members of $OPT(\varepsilon)$. ∎
A.5 Proof of Theorem 3.6
@@
A.6 Proof of Theorem 3.7
@@
Appendix B PMC Infeasibility Example